# Hopfield-Fenchel-Young Networks: A Unified Framework for Associative Memory Retrieval

### v2

@@ B.7 Proof of Proposition 14
We now bound $-\bm{r}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})$ in two possible ways. Using the Cauchy-Schwarz inequality, we have $-\bm{r}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})\leq\|\bm{X}\bm{r}\|\|\bm{y}_{i}-\bm{y}_{j}\|\leq\sigma_{\max}(\bm{X})D_{i}$, where $\sigma_{\max}(\bm{X})$ is the largest singular value of $\bm{X}$ (its spectral norm). On the other hand, denoting $R_{i}:=\max_{j}\|\bm{y}_{i}-\bm{y}_{j}\|_{1}$, we can also use Hölder’s inequality to obtain $-\bm{r}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})\leq\|\bm{X}\bm{r}\|_{\infty}\|\bm{y}_{i}-\bm{y}_{j}\|_{1}\leq MR_{i}$, where we used the fact that $\|\bm{X}\bm{r}\|_{\infty}=\max_{k}|\bm{x}_{k}^{\top}\bm{r}|\leq\|\bm{x}_{k}\|\|\bm{r}\|=M$. Combining the two inequalities, we obtain $\bm{q}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})\geq\Delta_{i}-\epsilon\min\{\sigma_{\max}(\bm{X})D_{i},MR_{i}\}$. Using the assumption that $\Delta_{i}\geq\frac{D_{i}^{2}}{2\beta}+\epsilon\min\{\sigma_{\max}(\bm{X})D_{i},MR_{i}\}$, we obtain $\bm{q}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})\geq\frac{D_{i}^{2}}{2\beta}$, which from the previous points ensures convergence to $\bm{X}^{\top}\bm{y}_{i}$ in one iteration. The result follows by noting that, since $\mathcal{Y}\subseteq\{0,1\}^{D}$, we have $R_{i}=D_{i}^{2}$.
@@ C.2 MNIST $K$-MIL
[…]

### v5

@@ B.7 Proof of Proposition 14
We now bound $\bm{r}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})$ in two possible ways. Using the Cauchy-Schwarz inequality, we have $\bm{r}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})\leq\|\bm{X}\bm{r}\|\|\bm{y}_{i}-\bm{y}_{j}\|\leq\sigma_{\max}(\bm{X})D_{i}$, where $\sigma_{\max}(\bm{X})$ is the largest singular value of $\bm{X}$ (its spectral norm). On the other hand, denoting $R_{i}:=\max_{j}\|\bm{y}_{i}-\bm{y}_{j}\|_{1}$, we can also use Hölder’s inequality to obtain $\bm{r}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})\leq\|\bm{X}\bm{r}\|_{\infty}\|\bm{y}_{i}-\bm{y}_{j}\|_{1}\leq MR_{i}$, where we used the fact that $\|\bm{X}\bm{r}\|_{\infty}=\max_{k}|\bm{x}_{k}^{\top}\bm{r}|\leq\|\bm{x}_{k}\|\|\bm{r}\|=M$. Combining the two inequalities, we obtain $\bm{q}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})\geq\Delta_{i}-\epsilon\min\{\sigma_{\max}(\bm{X})D_{i},MR_{i}\}$. Using the assumption that $\Delta_{i}\geq\frac{D_{i}^{2}}{2\beta}+\epsilon\min\{\sigma_{\max}(\bm{X})D_{i},MR_{i}\}$, we obtain $\bm{q}^{\top}\bm{X}^{\top}(\bm{y}_{i}-\bm{y}_{j})\geq\frac{D_{i}^{2}}{2\beta}$, which from the previous points ensures convergence to $\bm{X}^{\top}\bm{y}_{i}$ in one iteration. The result follows by noting that, since $\mathcal{Y}\subseteq\{0,1\}^{D}$, we have $R_{i}=D_{i}^{2}$.
@@ C.2 MNIST $K$-MIL
Table 7: Dataset sample details for the MNIST $K$-MIL experiment. The size $L_{i}$ of the $i\textsuperscript{th}$ bag is determined through $L_{i}=\max\{K,L_{i}^{\prime}\}$ where $L_{i}^{\prime}\sim\mathcal{N}(\mu,\sigma^{2})$. The number of positive instances in a bag is uniformly sampled between $K$ and $L_{i}$ for positive bags and between 0 and K1K-1 for negative bags.